-Put the PDA36A2/PDA10A2 (TRANS/REFL for 775 respectively) on the table. Powered them with a new strip above this side of the table.
-Strung up 30+ ft homemade BNC cables for these signals (labelled Etienne/Achane).
-Moved the "floating" NIR camera to the other side of the room, to be used for 775 Transmission for the time being. If cameras are needed come talk to me. We should get more NIR cameras I think.
-Made several space saving optics mounts/base changes to more easily fit on the very cramped FC table.
-Prepping initial alignment into the cavity (still with no mirrors). Want to clean the PLA 3D printed alignment tool Daniel made. Seems like Iso is okay for this. Tested it on a reject print, small amount of Iso on a chem wipe, wipe clean, dry it. Waited a few minutes. No visible deformation on the test part so cleaning the real one now.
-Cleaned 3D printed alignment tool. Spent a fair amount of time using the two 775 steering mirrors to align into the first two OFC2 mirror positions (again no super optics on the table yet). With the super optics (tomorrow?) in place I believe the 775 light should fall into alignment fairly fast, these align tools have worked well in the past.
-Same thing as above but the 1550 light. Input should be first pass aligned.
-Aligned the PD and floating camera to the 775 first pass Trans path.
-Roughly aligned the flip up mirror in the 1550 Trans path onto the 50:50 BS for used later.
-End of work particle count: 0/0/0
This is about everything I can think of to prep the second filter cavity. Will hopefully unbox 4 new super optics tomorrow.